# Moderating Harm: Benchmarking Large Language Models for Cyberbullying Detection in YouTube Comments

### v2

@@ Abstract
As online platforms grow, comment sections increasingly host harassment that undermines user experience and well-being. This study benchmarks three leading large language models, OpenAI GPT 4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus, on a corpus of 5 080 YouTube comments sampled from high abuse threads in gaming, lifestyle, food vlog, and music channels. The dataset comprises 1 334 harmful and 3 746 non-harmful messages in English, Arabic, and Indonesian, annotated independently by two reviewers with almost perfect agreement (Cohen’s $\kappa$ = 0.83). Using a unified prompt and deterministic settings, GPT-4.1 achieved the best overall balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Gemini flagged the highest share of harmful posts (recall = 0.875) but its precision fell to 0.767 due to frequent false positives. Claude delivered the highest precision at 0.920 and the lowest false positive rate of 0.022, yet its recall dropped to 0.720. Qualitative analysis showed that all three models struggle with sarcasm, coded insults, and mixed-language slang. These results underscore the need for moderation pipelines that combine complementary models, incorporate conversational context, and fine-tune underrepresented languages and implicit abuse. A de-identified version of the dataset and full prompts is publicly released to promote reproducibility and further progress in automated content moderation.
@@ I-B Motivation and Problem Statement
This is especially dangerous in cyberbullying, where the impact of a single overlooked comment can be deeply personal [12]. Unlike hate speech or spam, bullying is often subtle: expressed through mockery, group pressure, or repeated jabs that may seem harmless in isolation but accumulate over time. Many academic benchmarks fail to capture this complexity, relying on synthetic or crowdsourced datasets with clear-cut abuse [13, 14].
@@ I-C2 Quantify error patterns
Measure false positive and false negative rates and identify recurring misclassification themes.
@@ II-A AI Content Moderation and Toxic Language Detection
Automated content moderation has evolved from simple keyword-based filters to machine learning classifiers and, more recently, transformer-based models. Early systems relied on blacklists or logistic regression trained on hand-labeled corpora, but struggled with negation, sarcasm, and informal language. The advent of BERT [15], RoBERTa [16], and their derivatives enabled context-aware classification by capturing bidirectional dependencies and subtle phrasing nuances. These models improved performance on public benchmarks such as the Jigsaw Toxic Comment dataset and HateXplain [17], driving adoption in industry moderation pipelines.
@@ II-B Challenges in Detecting Implicit and Nuanced Abuse
Detecting implicit harm remains one of the most difficult challenges in AI-based moderation. Sarcasm, coded language, cultural slang, and subtextual insults frequently evade model detection, even in LLMs. For example, a phrase like “She’s always such a queen, right?” may be an insult depending on tone and context, but appears harmless in isolation. Studies in sarcasm detection [20], adversarial misspellings, irony detection, and tone modeling has attempted to address this gap, but progress remains limited [21].
@@ II-C Moderating Multilingual and Under-resourced Content
While English dominates most training corpora, abuse occurs in every language, and much of it is multilingual, transliterated, or mixed with emojis and slang. Studies consistently show significant drops in model accuracy when applied to languages such as Arabic [23] and Hindi [24], particularly when users employ spelling variation or obfuscation to evade filters [25]. Code-switching and transliteration further degrade detection performance, as models frequently miss common abusive patterns or misinterpret them out of context [26, 27].
@@ III-D Evaluated Models
III-E Prompt Design
All three models were queried with the identical, minimal instruction shown below. Temperature was fixed at 0.0 and the maximum response length at 5 tokens to enforce deterministic, one–token outputs.
@@ III-E Prompt Design
The prompt provides no label definitions beyond the binary mapping, compelling each model to infer an internal policy from the plain-language description while still accounting for nuance.
@@ IV-B Model-Level Classification Performance
GPT achieved the most balanced performance across the board. It reached an F1 score of 0.863, with high values for both precision at 0.887 and recall at 0.841. Claude stood out for its high precision, achieving 0.920, and maintained the lowest false positive rate at 0.022. However, this conservative stance came at the cost of recall, which was limited to 0.72. Gemini prioritized identifying as many harmful comments as possible, leading to the highest recall of 0.875, but its precision dropped to 0.767 and its false positive rate increased to 0.095.
IV-C Observations and Comparisons
@@ V-A False Positives and Surface-Level Flagging
In a small number of cases, only 12 out of 5 080, all three models (GPT, Claude, and Gemini) identified comments as harmful, while human reviewers did not. These examples typically involved informal phrasing, sarcasm, or emotionally expressive language. For instance, “Y’all are so toxic” was likely flagged due to tone, despite not targeting a specific individual. Another example, “Omg…321lbs! That’s so sad,” lacked context and may have been flagged merely for mentioning weight. These instances illustrate how all three models tend to rely on surface-level cues, such as slang or emotional keywords, without fully assessing intent or context.
@@ VI Future Work
The study surfaces several gaps that invite deeper investigation and practical follow-up.
VI-A Thread-Level and Multimodal Context
All three models evaluated a single comment in isolation. Future work should supply surrounding conversation threads, video transcripts, and visual cues so that the model can gauge sarcasm, piling-on, or context that flips a seemingly benign sentence into harassment.
VI-B Implicit and Coded Abuse
Many errors stemmed from sarcasm, memes, or coded references. A dedicated corpus that labels irony, metaphor, and insider jokes would let researchers test whether prompt engineering or specialised pre-training improves detection of covert hostility.
VI-C Ensemble and Risk-Tiered Pipelines
No single model led across every metric. A practical pipeline could run a high-recall model first, then pass flagged items through a high-precision filter before escalation to human review. Measuring cost, latency, and user impact of such cascades remains an open task.
VI-D User-Facing Feedback
Flagging comments is only one lever. Interventions that nudge commenters toward civility, or that explain why language is harmful, might reduce abuse at the source. Controlled trials comparing silent removal, warning prompts, and educational pop-ups would clarify which approach changes behaviour most effectively.
VI-E Long-Term Outcomes and Mental-Health Impact
Finally, future studies should trace whether improved moderation actually lowers anxiety, self-harm ideation, or community churn. Partnering with mental-health researchers and using de-identified longitudinal data can turn model gains into real-world well-being.
@@ VII Conclusion
This paper benchmarks OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus on 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. Each model was evaluated with the same prompt under deterministic settings. Gemini identified the largest share of harmful content with recall 0.875 but its precision fell to 0.767 due to frequent false positives. Claude achieved the highest precision 0.920 and the lowest false positive rate 0.022 yet its recall slipped to 0.720. GPT delivered the best overall balance with an F1 score 0.863, precision 0.887, and recall 0.841. Qualitative analysis showed that sarcasm, coded insults, and mixed-language slang remain persistent blind spots for all three models. These findings make clear that no single system meets every moderation requirement. Practical pipelines should combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and subtle forms of abuse. The fully de-identified dataset and detailed prompts are publicly released to support reproducible research and foster further improvements in automated content moderation.

### v3

@@ Abstract
As online platforms grow, comment sections increasingly host harassment that undermines user experience and well-being. This study benchmarks three state-of-the-art large language models: OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus, on a corpus of 5 080 YouTube comments drawn from high-abuse videos in gaming, lifestyle, food-vlog, and music channels. The dataset comprises 1 334 harmful and 3 746 non-harmful messages in English, Arabic, and Indonesian, annotated independently by two reviewers with almost perfect agreement (Cohen’s $\kappa=0.83$). Each model is evaluated in a strict zero-shot setting with an identical minimal prompt and deterministic decoding, giving a fair multi-language comparison without task-specific tuning. GPT-4.1 achieves the best balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Gemini flags the most harmful posts (recall = 0.875) but its precision falls to 0.767 because of frequent false positives. Claude attains the highest precision at 0.920 and the lowest false-positive rate of 0.022, yet its recall drops to 0.720. Qualitative analysis shows that all three models struggle with sarcasm, coded insults, and mixed-language slang. The findings highlight the need for moderation pipelines that combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and implicit abuse. A de-identified version of the dataset, full prompts, and model outputs is publicly released to promote reproducibility and further progress in automated content moderation.
@@ I-B Motivation and Problem Statement
This is especially dangerous in cyberbullying, where the impact of a single overlooked comment can be deeply personal [12]. Unlike hate speech or spam, bullying is often subtle: expressed through mockery, group pressure, or repeated jabs that may seem harmless in isolation but accumulate over time. Many academic benchmarks fail to capture this complexity, relying on synthetic or crowdsourced datasets with clear-cut abuse [13, 14]. Even widely used corpora such as OLID [15] and HateCheck [16] reveal the same gap: they consist of short, isolated text snippets and miss the multi-turn, emotionally layered exchanges common in YouTube threads.
@@ I-C2 Quantify error patterns
Measure false positive and false negative rates and identify recurring misclassifications themes.
@@ II-A AI Content Moderation and Toxic Language Detection
Automated content moderation has evolved from simple keyword-based filters to machine learning classifiers and, more recently, transformer-based models. Early systems relied on blacklists or logistic regression trained on hand-labeled corpora, but struggled with negation, sarcasm, and informal language. The advent of BERT [17], RoBERTa [18], and their derivatives enabled context-aware classification by capturing bidirectional dependencies and subtle phrasing nuances. These models improved performance on public benchmarks such as the Jigsaw Toxic Comment dataset and HateXplain [19], driving adoption in industry moderation pipelines. Mathew et al. [19] fine-tune BERT-base on HateXplain and report a macro-F1 of 0.674; a side-by-side comparison with our zero-shot LLM scores appears later in Table IV.
@@ II-B Challenges in Detecting Implicit and Nuanced Abuse
Detecting implicit harm remains one of the most difficult challenges in AI-based moderation. Sarcasm, coded language, cultural slang, and subtextual insults frequently evade model detection, even in LLMs. For example, a phrase like “She’s always such a queen, right?” may be an insult depending on tone and context, but appears harmless in isolation. Studies in sarcasm detection [22], adversarial misspellings, irony detection, and tone modeling have attempted to address this gap, but progress remains limited [23].
@@ II-C Moderating Multilingual and Under-resourced Content
Although English dominates most training corpora, abuse occurs in every language, and much of it is multilingual, transliterated, or mixed with emojis and slang. Studies consistently show significant drops in model accuracy when applied to languages such as Arabic [25] and Hindi [26], particularly when users employ spelling variation or obfuscation to evade filters [27]. Code-switching and transliteration further degrade detection performance, as models frequently miss common abusive patterns or misinterpret them out of context [28, 29].
@@ III-D Evaluated Models
III-E Prompt Design and Justification
All three models were queried with the same minimal prompt, using temperature = 0.0 to remove randomness and max_tokens = 10 to leave a small buffer for whitespace while still forcing a single-digit label. Locking these parameters serves three goals:
  1. 1.
    Fair cross-model comparison. One fixed prompt eliminates tuning bias and attributes performance differences solely to the models.
  2. 2.
    Deterministic, reproducible outputs. With temperature 0.0 and a short token cap, each comment always returns the same “0” or “1,” so any researcher can replicate the results with a single API call.
  3. 3.
    Scope control. The study provides a baseline snapshot of zero-shot LLM moderation on real-world data rather than an optimised prompt. Systematic prompt engineering is left to future work.
The fixed prompt used in every call is shown below.
@@ III-E Prompt Design and Justification
The prompt offers no label definitions beyond the binary mapping, requiring each model to infer its own policy from plain language while still accounting for nuance.
@@ IV-B Model-Level Classification Performance
GPT achieved the most balanced performance across the board. It reached an F1 score of 0.863, with high values for both precision at 0.887 and recall at 0.841. Claude stood out for its high precision, achieving 0.920, and maintained the lowest false positive rate at 0.022. However, this conservative stance came at the cost of recall, which was limited to 0.72. Gemini prioritized identifying as many harmful comments as possible, leading to the highest recall of 0.875, but its precision dropped to 0.767 and its false positive rate increased to 0.095. Representative misclassified comments are discussed in Section V-A.
IV-C Baseline comparison with a fine-tuned transformer
Mathew et al. [19] fine-tune BERT-base on the HateXplain corpus and report a macro-F1 of 0.674 (Table 6 of their paper). Table IV sets that published baseline beside the zero-shot scores obtained here. Because the corpora differ, these figures are not a head-to-head benchmark; the BERT row is included only as a widely cited transformer reference point.
TABLE IV: Macro-F1 for a classic fine-tuned baseline versus zero-shot LLMs
Model Dataset / Setting Macro-F1
BERT-base (fine-tuned) [19] HateXplain 0.674
GPT-4.1 (zero-shot, this work) YouTube 5 080 0.863
Gemini 1.5 Pro (zero-shot, this work) YouTube 5 080 0.818
Claude 3 Opus (zero-shot, this work) YouTube 5 080 0.808
IV-D Observations and Comparisons
@@ V-A False Positives and Surface-Level Flagging
In a small number of cases, only 12 out of 5 080, all three models (GPT, Claude, and Gemini) identified comments as harmful, while human reviewers did not. These examples typically involved informal phrasing, sarcasm, or emotionally expressive language. For instance, the comment “Y’all are so toxic” was likely flagged due to tone, despite not targeting a specific individual. Another example, “Omg…321lbs! That’s so sad,” lacked context and may have been flagged merely for mentioning weight. These instances illustrate how all three models tend to rely on surface-level cues, such as slang or emotional keywords, without fully assessing intent or context.
@@ VI Future Work
The open issues are ranked below from highest near-term impact to longer-term research challenges.
  1. 1.
    Thread-level and multimodal context. Adding preceding turns, video transcripts, or image cues is the fastest way to cut many sarcasm and piling-on errors found in this study.
  2. 2.
    Ensemble and risk-tiered pipelines. A two-stage approach (first a high-recall filter, then a high-precision check) can be deployed with existing APIs. Measuring latency, cost, and user impact belongs in the next round of experiments.
  3. 3.
    Implicit and coded abuse corpora. New datasets that label irony, metaphor, and insider slang will let researchers test prompt tweaks and specialised pre-training for covert hostility. This requires a dedicated annotation effort.
  4. 4.
    User-facing feedback loops. Controlled trials comparing silent removal, warning prompts, and educational pop-ups could clarify which intervention nudges commenters toward civility most effectively.
  5. 5.
    Long-term mental-health outcomes. Ultimately we need evidence that better moderation lowers anxiety, self-harm ideation, or community churn. Partnering with mental-health researchers and analysing de-identified longitudinal data are goals for later phases.
@@ VII Conclusion
This paper benchmarks OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus on 5 080 YouTube comments drawn from high-abuse threads in gaming, lifestyle, food vlog, and music channels. Each model was evaluated with the same prompt under deterministic settings. Gemini identified the largest share of harmful content, achieving recall of 0.875, yet its precision dropped to 0.767 because of frequent false positives. Claude reached precision of 0.920 and the lowest false-positive rate of 0.022, although its recall fell to 0.720. GPT-4.1 delivered the best overall balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Qualitative analysis showed that sarcasm, coded insults, and mixed-language slang remain persistent blind spots for all three models. These findings make clear that no single system meets every moderation requirement. Practical pipelines should combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and subtle forms of abuse.
For deployment, platforms can map each model’s strengths to specific risk tiers. A site that values maximum coverage can run Gemini as the first-pass filter and then route its flags to Claude or a human queue to reduce false alarms. Low-latency chat services may prefer using GPT alone for a balanced trade-off between recall and precision. Logging each disagreement and feeding it into periodic retraining supports continuous improvement, auditability, and fairness, contributing to AI-safety goals that limit user exposure to harmful content without over-silencing benign speech.
The fully de-identified dataset, prompt text, and model outputs are openly released to support reproducible research and to foster further progress in automated content moderation.
@@ VIII Limitations
To address the cultural-bias risk noted in the Limitations section, future releases will include reviews from annotators outside the authors’ demographic group and will make the labeling guide public so external researchers can audit or challenge specific decisions.